OpenStudy (anonymous):

Use the substitution method to solve the system of equations. 4x – y = 4 x + 4y = 18 (4, 2) (1, 4) (4, 1) (2, 4) Use the substitution method to solve the system of equations. 4x – y = 4 x + 4y = 18 (4, 2) (1, 4) (4, 1) (2, 4) @Mathematics

hero (hero):

4x-4 = y (18-x)/4 = y y = y 4(4x-4) = 18 -x 16x-16 = 18 - x 17x = 34 x = 2 y = 4